OpenStudy (anonymous):

You exert 375 J to move a 45-kg object 5 m. How much force is required? A. 1,875 N B. 75 N C. 420 N D. 225 N

OpenStudy (anonymous):

I didnt even do this lesson >.<

OpenStudy (amistre64):

w = force*distance

OpenStudy (amistre64):

dont recall the newton joule kg-m stuff tho

OpenStudy (anonymous):

Hmm alright (:

OpenStudy (amistre64):

but... force = w/d adn force = m*a ...

OpenStudy (anonymous):

Okay :D :D

OpenStudy (amistre64):

75 is my best uneducated guess...

OpenStudy (anonymous):

Okay (: Btw.... you have effing 1098 medals!!!!!!!! Jeezus amistre :D